OpenStudy (bananas1234):

Express the following quotient in scientific notation. (3.472 × 10^-12) ÷ (5.6 × 10^-3)

OpenStudy (anonymous):

You plot those two terms in your scientific calculator and it shall give you in scientific notation form:\[\frac{ 3.472*10^{-12} }{ 5.6*10^{-3} }=6.2*10^{-10}\]

OpenStudy (anonymous):

@UnkleRhaukus it is division, not multiplication.

OpenStudy (bananas1234):

Thank you, i think i understand that a bit better now.